OpenStudy (anonymous):

sqrt{9x+81} =x+5 sqrt{9x+81} =x+5 @Mathematics

OpenStudy (anonymous):

\[\sqrt{9x+81 } = x+5\]

OpenStudy (agreene):

x=7

OpenStudy (anonymous):

i see what i did wrong..I added wrong

OpenStudy (anonymous):

x=7